# Utility-Diversity Aware Online Batch Selection for LLM Supervised Fine-tuning

### v1

@@ B.1 Parameter Analysis
We evaluate the influence of the projected dimensions $d_{1}$, $d_{2}$, and the memory buffer size $M$ on both model performance and resource usage. As shown in Figures 5 and 6, increasing the projected dimensions and memory buffer consistently improves accuracy, which gradually saturates once $d_{1}$, $d_{2}$, and $M$ become sufficiently large. The trends of $d_{1}$ and $d_{2}$ are consistent with the Johnson-Lindenstrauss lemma (Johnson et al., 1984; Ailon & Chazelle, 2006; Jin et al., 2021; Ailon & Chazelle, 2009; Tropp, 2011), while a sufficiently large $M$ leads to performance saturation, indicating that the stored representations are adequate to capture global sample diversity. The corresponding increases in memory usage and computation time are minor, and arise only when computing inter-sample diversity.
@@ C.2 Optimization Analysis using Adam
When using the Adam (Kingma & Ba, 2014) as optimizer, the parameter update at step $t$ is
@@ C.3 Distance Preservation under Low-dimensional Projection
Let $\bm{u}_{1},\dots,\bm{u}_{n}\in\mathbb{R}^{NV}$ denote $n$ vectors of interest. Consider all pairwise differences $\bm{u}_{i}-\bm{u}_{j}$, $1\leq i<j\leq n$; there are at most n(n1)/2n(n-1)/2 such vectors. By standard results on Fast JL transforms (SRFT/FJLT) (Ailon & Chazelle, 2006; Jin et al., 2021; Ailon & Chazelle, 2009; Tropp, 2011), for any $0<\epsilon<1$ and a fixed set of $n$ vectors in $\mathbb{R}^{NV}$, a random SRFT $\bm{\Gamma}\in\mathbb{R}^{d_{1}d_{2}\times NV}$ preserves the $\ell_{2}$ norms up to $(1\pm\epsilon)$ with high probability provided
@@ D.1 Datasets
  • MMLU Hendrycks et al. (2021) is a benchmark designed to assess general knowledge understanding, consisting of multiple-choice questions from various branches of knowledge. It covers 57 tasks including elementary mathematics, US history, computer science, law, and more.
  • ScienceQA Lu et al. (2022) is a multimodal science question answering dataset collected from elementary and high school science curricula, including multiple-choice problems that align with California Common Core Content Standards. The questions in dataset are sourced from open resources managed by IXL Learning, an online learning platform curated by experts in the field of K-12 education. We extract textual parts within it following Li et al. (2024).
  • GSM8K Cobbe et al. (2021) is a dataset of 8.5K high quality linguistically diverse grade school math word problems, each requiring multi-step arithmetic reasoning. Answers are provided in a step-by-step explanation format.
  • CodeAlpaca-20k Chaudhary (2023) is an instruction-following dataset consisting of 20k examples generated to align code-related prompts with helpful outputs. It is synthetically constructed to enhance code instruction tuning.
  • HumanEval Chen et al. (2021) is a code generation benchmark including 164 Python programming problems. The dataset was handwritten to ensure not to be included in the training set of code generation models.
@@
MMLU Hendrycks et al. (2021) $99842$ $14042$ Multiple Choice
ScienceQA Lu et al. (2022) $12726$ $4241$ Multiple Choice
GSM8K Cobbe et al. (2021) $7473$ $1319$ Math Problems
CodeAlpaca-20k Chaudhary (2023) $20022$ - Code Instruction
HumanEval Chen et al. (2021) - $164$ Code Generation
@@ Appendix E Limitations and Future Work
One potential limitation of our framework lies in the computation of the nuclear norm. While it can be obtained directly from the singular values of the logits matrix without incurring additional expensive backpropagation, the SVD of large matrices may still introduce non-negligible overhead in practice, particularly for long sequences with large vocabulary sizes. We also explored approximate methods such as randomized SVD to accelerate the computation, but observed that the loss in precision often degrades the overall performance.
@@
An important future direction is to explore more efficient and accurate estimators of the nuclear norm. Potential avenues include exploiting structured low-rank approximations, leveraging iterative or block-wise SVD techniques, or designing surrogate scoring functions that preserve the optimization and diversity signals of the nuclear norm while being cheaper to compute. Such improvements would further enhance the scalability of our framework to larger models and longer input sequences. Furthermore, techniques in uncertainty estimation or active learning may be other promising directions for online batch selection (Ren et al., 2021; Loquercio et al., 2020; Peng et al., 2024).

### v2

@@ B.1 Parameter Analysis
We evaluate the influence of the projected dimensions $d_{1}$, $d_{2}$, and the memory buffer size $M$ on both model performance and resource usage. As shown in Figures 6 and 7, increasing the projected dimensions and memory buffer consistently improves accuracy, which gradually saturates once $d_{1}$, $d_{2}$, and $M$ become sufficiently large. The trends of $d_{1}$ and $d_{2}$ are consistent with the Johnson-Lindenstrauss lemma (johnson1984extensions; ailon2006approximate; jin2021faster; ailon2009fast; tropp2011improved), while a sufficiently large $M$ leads to performance saturation, indicating that the stored representations are adequate to capture global sample diversity. The corresponding increases in memory usage and computation time are minor, and arise only when computing inter-sample diversity.
@@
B.2 Additional Ablation Studies on Buffer Update Policies, Distance Aggregation Strategies, and Random Matrix Constructions
For buffer update policies, we compare our default FIFO strategy with two alternatives: reservoir sampling (random replacement within the buffer) and class-aware sampling (constructing class prototypes via K-means). Experiments on the MMLU benchmark using Qwen-2.5-7B (Table 7) show that all three policies yield comparable performance, with no notable differences observed.
For distance aggregation strategies, we evaluate our average-distance scheme against two variants—farthest-$k$ and soft-min—again on MMLU with Qwen-2.5-7B (Table 8). The results similarly indicate no significant performance differences among these methods.
In UDS, we employ an SRFT-style construction for the bidirectional random projection matrix. To assess its necessity, we compare it against two alternative constructions: Sparse JL and CountSketch-style tensor sketches. Experiments on MMLU with Qwen-2.5-7B (Table 9) show that the CountSketch-style transform achieves performance nearly identical to the SRFT-style transform, whereas Sparse JL performs significantly worse.
The inferior performance of Sparse JL stems from its failure to satisfy the JL lemma under the bidirectional formulation, similar to the Gaussian-style construction discussed in Section 3.2. The issue arises because the Kronecker product disrupts the independence structure of the random variables, thereby violating the distance-preserving requirements of the JL lemma. In contrast, both SRFT-style and CountSketch-style constructions continue to satisfy the JL lemma when used bidirectionally.
From an efficiency perspective, decomposing the full random projection into a bidirectional one already greatly reduces memory usage and computation cost, making these overheads negligible relative to other components of the algorithm. Consequently, as long as a construction satisfies the JL lemma in the bidirectional setting, it is unnecessary to further distinguish which one is marginally more efficient. The SRFT-style construction is therefore sufficient for our purposes, and is what we adopt in our main experiments.
Table 7: Ablation Studies on Buffer Update Policies. We report average accuracy on MMLU using Qwen-2.5-7B.
FIFO Reservoir Sampling Class-aware Sampling
[…]
Table 8: Ablation Studies on Distance Aggregation Policies. We report average accuracy on MMLU using Qwen-2.5-7B.
Average Distance Farthest-k Soft-min
[…]
Table 9: Ablation Studies on Different Construction of the Random Matrix. We report average accuracy on MMLU using Qwen-2.5-7B.
SRFT-style Sparse JL CountSketch-style
[…]
B.3 Experiments on Full SFT, Larger Batch Sizes, and Instruction-Tuned Models
To further demonstrate the scalability of UDS, we provide additional results under three settings: larger batch sizes (mini-batch size $B=16$ in Table 10), full-parameter finetuning with Qwen-2.5-7B (Table 11), and instruction-tuned models using Qwen-2.5-7B-Instruct (Table 12). Across all scenarios, UDS consistently outperforms the corresponding baselines, strengthening the robustness and generalization capability of the method.
Table 10: Accuracy Comparison for Different Online Batch Selection Methods with Larger Batch Size. We evaluate various methods across four benchmarks: MMLU, ScienceQA, GSM8K, and HumanEval. $\bar{A}$ denotes average accuracy or Pass@1 reported in percentage ($\%$). The best results of $\bar{A}$ are highlighted in bold.
Model Method MMLU ScienceQA GSM8K HumanEval
[…]
Table 11: Accuracy Comparison for Different Online Batch Selection Methods with Full Finetuning. We evaluate various methods across four benchmarks: MMLU, ScienceQA, GSM8K, and HumanEval. $\bar{A}$ denotes average accuracy or Pass@1 reported in percentage ($\%$). The best results of $\bar{A}$ are highlighted in bold.
Model Method MMLU ScienceQA GSM8K HumanEval
[…]
Table 12: Accuracy Comparison for Different Online Batch Selection Methods on Instruction Tuning Model. We evaluate various methods across four benchmarks: MMLU, ScienceQA, GSM8K, and HumanEval. $\bar{A}$ denotes average accuracy or Pass@1 reported in percentage ($\%$). The best results of $\bar{A}$ are highlighted in bold.
Model Method MMLU ScienceQA GSM8K HumanEval
[…]
B.4 Additional Experiments on Tasks Requiring Long-Context Reasoning
In all previous experiments, we set the maximum sequence length to 512. To assess whether UDS can effectively capture utility and diversity for training samples that require long-context reasoning, we further conduct experiments on the MATH dataset (hendrycks2021measuring), using a maximum sequence length of 2048. We train on the MATH training split and evaluate on its test split. As shown in Table 13, UDS continues to deliver strong performance compared with other baselines, even under this more challenging long-context setting.
Table 13: Additional Studies on MATH dataset with Longer CoT Reasoning. We report average accuracy using Qwen-2.5-7B.
[…]
B.5 Additional Experiments on OOD Evaluation
To further assess the robustness of UDS, we conduct out-of-distribution (OOD) evaluations using GSM8K as a case study. Specifically, we train on the GSM8K training set and evaluate on its in-distribution test set as well as two OOD benchmarks: MATH500 and AMC23 (lightman2023lets). As shown in Table 14, UDS continues to exhibit strong performance across both in-distribution and OOD settings, demonstrating its robustness to distribution shifts.
Table 14: Accuracy Comparison for Different Online Batch Selection Methods on OOD Datasets. We evaluate various methods across three benchmarks: GSM8K (in-distribution), MATH500, and AMC23 (out-of-distribution). $\bar{A}$ denotes average accuracy or Pass@1 reported in percentage ($\%$). The best results of $\bar{A}$ are highlighted in bold.
Model Method GSM8K MATH500 AMC23
[…]
B.6 Strong Correlation between the Frobenius Norm of the Logits Matrix and Sample Utility
Similar to Figures 2 and 3, we also observe a strong empirical correlation between the Frobenius norm of the logits matrix, $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$, and the change in loss, $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$, as illustrated in Figure 9.
Figure 9: Strong correlation between $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$ during the training process. Each point represents the correlation coefficient calculated from $B$ pairs of values within a single batch. We extract the result using Qwen-2.5-7B on MMLU.
B.7 Comparison with Offline Data Selection Methods
We additionally include a representative offline data selection method, FisherSFT (deb2025fishersft), which uses Fisher information gain to select a training subset, in our comparison. We conduct experiments on all four datasets using Qwen-2.5-7B as the backbone model. Since FisherSFT does not operate in the same manner as online batch selection methods, it is infeasible to fairly compare throughput, so we primarily compare the final accuracy after training. Both FisherSFT and UDS select the same fraction of data, and all other training details remain the same. The results in Table 15 show that UDS also achieves better performance than FisherSFT.
Table 15: Accuracy Comparison with Offline Data Selection Strategy FisherSFT. We report average accuracy using Qwen-2.5-7B.
Method MMLU ScienceQA GSM8K HumanEval
[…]
@@ C.2 Optimization Analysis using Adam
When using the Adam (kingma2014adam) as optimizer, the parameter update at step $t$ is
@@ C.3 Distance Preservation under Low-dimensional Projection
Let $\bm{u}_{1},\dots,\bm{u}_{n}\in\mathbb{R}^{NV}$ denote $n$ vectors of interest. Consider all pairwise differences $\bm{u}_{i}-\bm{u}_{j}$, $1\leq i<j\leq n$; there are at most n(n1)/2n(n-1)/2 such vectors. By standard results on Fast JL transforms (SRFT/FJLT) (ailon2006approximate; jin2021faster; ailon2009fast; tropp2011improved), for any $0<\epsilon<1$ and a fixed set of $n$ vectors in $\mathbb{R}^{NV}$, a random SRFT $\bm{\Gamma}\in\mathbb{R}^{d_{1}d_{2}\times NV}$ preserves the $\ell_{2}$ norms up to $(1\pm\epsilon)$ with high probability provided
@@ D.1 Datasets
  • MMLU (hendryckstest2021) is a benchmark designed to assess general knowledge understanding, consisting of multiple-choice questions from various branches of knowledge. It covers 57 tasks including elementary mathematics, US history, computer science, law, and more.
  • ScienceQA (lu2022learn) is a multimodal science question answering dataset collected from elementary and high school science curricula, including multiple-choice problems that align with California Common Core Content Standards. The questions in dataset are sourced from open resources managed by IXL Learning, an online learning platform curated by experts in the field of K-12 education. We extract textual parts within it following li2024mixlora.
  • GSM8K (cobbe2021gsm8k) is a dataset of 8.5K high quality linguistically diverse grade school math word problems, each requiring multi-step arithmetic reasoning. Answers are provided in a step-by-step explanation format.
  • CodeAlpaca-20k (codealpaca) is an instruction-following dataset consisting of 20k examples generated to align code-related prompts with helpful outputs. It is synthetically constructed to enhance code instruction tuning.
  • HumanEval (chen2021evaluating) is a code generation benchmark including 164 Python programming problems. The dataset was handwritten to ensure not to be included in the training set of code generation models.
@@
MMLU (hendryckstest2021) $99842$ $14042$ Multiple Choice
ScienceQA (lu2022learn) $12726$ $4241$ Multiple Choice
GSM8K (cobbe2021gsm8k) $7473$ $1319$ Math Problems
CodeAlpaca-20k (codealpaca) $20022$ - Code Instruction
HumanEval (chen2021evaluating) - $164$ Code Generation
@@ Appendix E Limitations and Future Work
One potential limitation of our framework lies in the computation of the nuclear norm. While it can be obtained directly from the singular values of the logits matrix without incurring additional expensive backpropagation, the SVD of large matrices may still introduce non-negligible overhead in practice, particularly for long sequences with large vocabulary sizes. We also explored approximate methods such as randomized SVD to accelerate the computation, but observed that the loss in precision often degrades the overall performance. Another limitation is the difficulty of providing a rigorous theoretical proof for the linear correlation between the nuclear norm of the logits matrix and the resulting loss reduction or effective matrix rank.
@@
An important future direction is to explore more efficient and accurate estimators of the nuclear norm. Potential avenues include exploiting structured low-rank approximations, leveraging iterative or block-wise SVD techniques, or designing surrogate scoring functions that preserve the optimization and diversity signals of the nuclear norm while being cheaper to compute. Such improvements would further enhance the scalability of our framework to larger models and longer input sequences. Furthermore, techniques in uncertainty estimation or active learning may be other promising directions for online batch selection (ren2021survey; loquercio2020general; peng2024energy).